HelpMePlz:

Which of the following ordered pairs lies on the graph of h(x) = -2x²? (1, -4) (1, -2) (1, 4)

Champion:

(2,3) means x=2, y=3. To see if (2,3) is on the graph of h(x) = -2x² substitute x=2, and see if the result y=h(2)=-2(2)² equals 3. Since the result y=h(2)=-2(2)²=8 does not equal 3, we conclude that (2,3) is not on the graph of h(x). Do this similarly for all options to find the point on the graph.

HelpMePlz:

ok thanks